Project

General

Profile

12132013 Minutes

UMn....

Report from Siva....
The NuMI trigger rate is little down the previous and is fluctuating.
dcm-2-05-02 is functioning since morning(warm, 350V).
A few dcm's 2-03{10,12}, dcm-2-5-{3,5} are noisy since last night.
These issues I want to talk about today in the meeting.

Operationally: Running All APDs cold on diblocks 1,2,3 cold
Running dcm 4-1 and 5-2 at full gain warm now.

Dom - FTS is a little bit behind. (Maybe 5 runs). We will monitor.